### v3

# Isoparametric functions on Finsler space forms

Abstract: In this paper, we prove that transnormal functions are isoparametric functions on Finsler space forms $(N(c),F)$ under certain conditions, which generalize Theorem B given by Q. M. Wang in Riemannian case. Next, we discuss the relationship between umbilic hypersurfaces and isoparametric hypersurfaces in $(N(c),F)$. Further more, we construct a locally isoparametric function by a distance function and give a global isoparametric function on a standard Finsler sphere.
Key words: Finsler space form; Transnormal function; Isoparametric function; Umbilic; Distance function.
Mathematics Subject Classification (2010): 53C20, 53C60, 53C42.
1 Introduction
The study on isoparametric theory has a long history. A smooth function $f$ defined on a connected complete Riemannian manifold $(N,g)$ is called an isoparametric function if there exists a smooth function $\bar{a}(t)$ and a continuous function $\bar{b}(t)$ satisfying
$\left\{\begin{aligned} &\|\nabla f\|^{2}=\bar{a}(f),\\ &\Delta f=\bar{b}(f).\end{aligned}\right$. (1.1)
E. Cartan began to study isoparametric hypersurfaces in real space forms with constant sectional curvature $c$ systematically [1]. Then many mathematicians started working on isoparametric functions and the classification of isoparametric hypersurfaces in real space forms [2, 3, 4]. If $f$ only satisfies the first equation of (1.1), $f$ is called a transnormal function.
The fundamental structural result in a general Riemannian manifold $N$ was given by Q. M. Wang [15]. It shows that the focal varieties of a transnormal function $f$, which is the preimage if the maximum (resp. minimum), are smooth submanifolds of $N$ and each regular level set of $f$ is a tube over either of the focal varieties. Particularly, he obtained the following startling theorems,
Theorem A.
[15] Transnormal functions on $\mathbb{S}^{n}$ and $\mathbb{E}^{n}$ are isoparametric.
However, no proof was given in [15]. Later,  [14] supplied the proof of Theorem A and constructed an example to point out there is an essential difference between transnormal functions and isoparametric functions on hyperbolic spaces.
In Finsler geometry, [9] introduced the concept of isoparametric functions and transnormal functions. Let $(N,F)$ be an $n$-dimensional Finsler manifold. A $C^{1}$ non-constant function $f$ on $N$ is called isoparametric if it is smooth on $N_{f}=\{x\in N\mid df(x)\neq 0\}$ and there exist a smooth function $a(t)$ and a continuous function $b(t)$ defined on $J=f(N_{f})$ such that
$\left\{\begin{aligned} &F^{2}(\nabla f)=a(f),\\ &\Delta f=b(f).\end{aligned}\right$. (1.2)
Each regular level hypersurface $M_{t}=f^{-1}(t)$ is called an isoparametric hypersurface. The function satisfies the first equation in (1.2) is called a transnormal function.
As in Riemannian geometry, we call a complete and simply connected Finsler manifold with constant flag curvature a Finsler space form. There are many kinds of special Finsler space forms, such as Minkowski spaces (with zero flag curvature) and standard Finsler spheres (with positive flag curvature and vanishing $\mathbf{S}_{BH}$-curvature) [19]. If a Finsler manifold is only forward (resp. backward) complete, we call it forward (resp. backward) Finsler space form[10, 11, 6] classified isoparametric hypersurfaces in some kinds of Finsler space forms. [12] studied isoparametric hypersurfaces in Finsler space forms with vanishing reversible torsion.
The fundamental structural result in [15] was generalized to Finsler geometry in [5]. Based on the structural result given in [5], in this paper, we try to generalize Theorem A to Finsler geometry. Unlike the Riemannian case, the geodesics are not necessarily reversible and the regularity of $f$, $a(t)$ and $b(t)$ in (1.2) are very poor, which causes more difficulties for the research.
For a connected Finsler manifold $(N,F)$, set $\overline{J}=f(N)$. Denote the focal variety of $f$ by $M_{\pm}$, where $M_{+}$ (resp. $M_{-}$) is the set that $f$ attains its global maximum (resp. minimum) value. Let $f$ be a $C^{2}$ transnormal function satisfying $F^{2}(\nabla f)=a(f)$, where $a(t)\neq 0$ for $t\in J$ and $a^{\prime}(t)\neq 0$ for $t\in\partial\overline{J}$. We called such $f$ an appropriate transnormal function. Let $g$ be the number of distinct principal curvatures for an isoparametric hypersurface. Then we have
Theorem 1.1.
If $f$ is an appropriate transnormal function and f-f is a transnormal function on a Finsler space form $(N(c),F)$, then $f$ is isoparametric when $c\geq 0$, or $c<0$ and all principal curvatures $\kappa_{i}$ of a regular level hypersurface satisfying $|\kappa_{i}|\geq\sqrt{-c}$, $1\leq i\leq g$.
Remark 1.2.
Let $M_{\pm}$ be submanifolds and $f$ has no critical value in $J$. Then the condition in Theorem 1.1 that $f$ is an appropriate transnormal function can be weaken to $f$ is a transnormal function.
Theorem 1.3.
Let $(N(c),F)$ be a connected forward Finsler space form and $f$ be a transnormal function on $N$. If some level hypersurface of $f$ is umbilic, then $f$ is isoparametric.
Theorem 1.4.
Let $(N(c),F)$ be a forward Finsler space form and $r(x)=d_{F}(p,x)$ be a distance function from any $p\in N$. Then $f=\frac{1}{2}r^{2}(x)$ is a locally isoparametric function with $g=1$ on the cut-domain of $p$. Particularly, $f=-\cos(\sqrt{c}r)$ is a global isoparametric function on a standard Finsler sphere.
2 Preliminaries
In this section, we will give some definitions and lemmas that will be used in the proof of our main results. Here and from now on, we will use the following convention of index ranges unless other stated,
$1\leq i,j,\dots\leq n,\ \ \ \ 1\leq a,b,\dots\leq n-1.$
2.1 Finsler manifolds
Let $(N,F)$ be an $n$-dimensional manifold and let $TN=\cup_{x\in N}T_{x}N$ be the tangent bundle over $N$ with local coordinates $(x,y)$, where $x=(x^{i})$ and $y=y^{i}\frac{\partial}{\partial x^{i}}$. The fundamental form $g$ of $(N,F)$ is given by
$g=g_{ij}(x,y)dx^{i}\otimes dx^{j},~~~~~~~g_{ij}(x,y)=\frac{1}{2}[F^{2}]_{y^{i}y^{j}}.$
The geodesics $\gamma=\gamma(t)$ of $F$ in local coordinates $(x^{i})$ are characterized by
$\frac{\textmd{d}^{2}x^{i}}{\textmd{d}t^{2}}+2G^{i}(x,\frac{\textmd{d}x}{\textmd{d}t})=0,$ (2.1)
where $(x^{i}(t))$ are the coordinates of $\gamma(t)$ and $G^{i}=G^{i}(x,y)$ are defined by
$G^{i}=\frac{g^{il}}{4}\{[F^{2}]_{x^{k}y^{l}}y^{k}-[F^{2}]_{x^{l}}\},$
which are called the geodesic coefficients.
Let $\pi:TN\rightarrow N$ be the natural projection. There exists the unique Berwald connection $\nabla$ with $\nabla\frac{\partial}{\partial x^{i}}=\Gamma^{i}_{jk}dx^{k}\otimes\frac{\partial}{\partial x^{j}}$ on the pull-back bundle $\pi^{\ast}TN$, which satisfies $\Gamma^{i}_{jk}=\frac{\partial G^{i}}{\partial y^{j}\partial y^{k}}=\frac{\partial N^{i}_{j}}{\partial y^{k}}$. Let $\frac{\delta}{\delta x^{i}}:=\frac{\partial}{\partial x^{i}}-N^{j}_{i}\frac{\partial}{\partial y^{j}}$. The Riemannian curvature tensor is defined by
$R^{i}_{k}:=y^{j}R^{i}_{j~kl}y^{l},\ \ \ \ R_{jk}:=g_{ij}R^{i}_{k}=y^{i}R_{ijkl}y^{l},$
where
$R^{i}_{j~kl}=\frac{\delta\Gamma^{i}_{jl}}{\delta x^{k}}-\frac{\delta\Gamma^{i}_{jk}}{\delta x^{l}}+\Gamma^{i}_{km}\Gamma^{m}_{jl}-\Gamma^{i}_{lm}\Gamma^{m}_{jk}.$
The non-Riemannian curvature tensor is defined by
$P^{i}_{j~kl}=-F\frac{\partial\Gamma^{i}_{jk}}{\partial y^{l}}=-F(G)^{i}_{y^{j}y^{k}y^{l}}.$ (2.2)
For a vector $v=v^{i}\frac{\partial}{\partial x^{i}}$ satisfying $g_{ij}v^{i}v^{j}=1$ and $g_{ij}y^{i}v^{j}=0$, the flag curvature of $(N,F)$ is defined by
$K(y,v):=F^{-2}R_{jk}v^{j}v^{k}.$
For $X=X^{i}\frac{\partial}{\partial x^{i}}\in\Gamma(TN)$, the covariant derivative of $X$ along $v=v^{i}\frac{\partial}{\partial x^{i}}\in T_{x}N$ with respect to a reference vector $w\in T_{x}N\setminus\{0\}$ is defined by
$\nabla^{w}_{v}X(x)=\{v^{j}\frac{\partial X^{i}}{\partial x^{j}}(x)+\Gamma^{i}_{jk}(w)v^{j}X^{k}(x)\}\frac{\partial}{\partial x^{i}}.$ (2.3)
Then (2.1) can be expressed by $\nabla^{\dot{\gamma}}_{\dot{\gamma}}\dot{\gamma}\equiv 0$. If a vector field $J(s)$ along $\gamma(s)$ satisfies
$\nabla_{\dot{\gamma}}^{\dot{\gamma}}\nabla_{\dot{\gamma}}^{\dot{\gamma}}J+\textbf{R}_{\dot{\gamma}}(J)=0,$ (2.4)
then $J(s)$ is called the Jacobi field along $\gamma(s)$, where $\textbf{R}_{y}:=R^{i}_{k}(y)\frac{\partial}{\partial x^{i}}\otimes dx^{k}$. Let $(N,F)$ be a forward complete Finsler space. For $y\in S_{x}N=\{y\in TN\setminus\{0\}\big{|}F(y)=1\}$, define $\textbf{c}_{y}>0$ to be the first number $d>0$ such that there is a Jacobi field $J(s)$ along $\gamma(s)=\exp_{x}(sy)$, $0<s<d$, satisfying $J(0)=0=J(d)$. $\textbf{c}_{y}$ is called the conjugate value of $y$. From Cartan-Hadamard Theorem, if $K\leq 0$, $\textbf{c}_{y}=\infty$. Let $\textbf{i}_{y}$ be the supremum of $r>0$ such that $\gamma(s)$ is minimizing on $[0,r]$. We call $\mathcal{D}_{x}:=\{\exp_{x}(ty),0\leq t<\textbf{i}_{y},y\in S_{x}N\}$ the cut-domain of $x$ [17].
Let ${\mathcal{L}}:TN\rightarrow T^{\ast}N$ denote the Legendre transformation, satisfying ${\mathcal{L}}(\lambda y)=\lambda{\mathcal{L}}(y)$ for all $\lambda>0,~y\in TN$. Moreover,
$\mathcal{L}^{-1}(\xi)=F^{*}(\xi)[F^{*}]_{\xi^{i}}(\xi)\frac{\partial}{\partial x^{i}},\ \ \forall\xi\in T^{*}N\setminus\{0\},\ \ \mathcal{L}^{-1}(0)=0,$
where $F^{*}$ is the dual metric of $F$. For a smooth function $f:N\rightarrow\mathbb{R}$, the gradient vector of $f$ at $x$ is defined as $\nabla f(x)={\mathcal{L}}^{-1}(df(x))\in T_{x}N$. Set $\nabla^{2}f(x)=\nabla^{\nabla f}(\nabla f)(x)$ for $x\in N_{f}$. The Hessian of $f$ is defined by
$\textrm{Hes}f(X,Y):=g_{\nabla f}(\nabla^{2}f(X),Y)=X(Yf)-(\nabla^{\nabla f}_{X}Y)f,$ (2.5)
where $X,Y\in T_{x}N$, $\nabla^{2}f(X)=\nabla^{\nabla f}_{X}\nabla f\in T_{x}N$ for $x\in N_{f}$. The Laplacian of $f$ can defined by
$\hat{\Delta}f=\textmd{tr}_{g_{{}_{\nabla f}}}(\nabla^{2}f).$ (2.6)
And Laplacian of $f$ with respect to the volume form $d\mu=\sigma(x)dx=\sigma(x)dx^{1}\wedge dx^{2}\wedge\cdots\wedge dx^{n}$ can be represented as
$\Delta_{\sigma}f=\textmd{div}_{\sigma}(\nabla f)=\frac{1}{\sigma}\frac{\partial}{\partial x^{i}}(\sigma g^{ij}(\nabla f)f_{j})=\hat{\Delta}f-\textbf{S}(\nabla f),$ (2.7)
where
$\textbf{S}(x,y)=\frac{\partial G^{i}}{\partial y^{i}}-y^{i}\frac{\partial}{\partial x^{i}}(\ln\sigma(x))$
is the $\mathbf{S}$-curvature [17].
2.2 Anisotropic submanifolds
Let $(N,F)$ be an $n$-dimensional Finsler manifold and $\phi:M\to(N,F)$ be an (n1)(n-1)-dimensional immersion. Set $(x,y)\in TN$ with local coordinate $(x^{i},y^{i})$ and $(u,v)\in TM$ with local coordinate $(u^{a},v^{a})$, we have
$x^{i}=\phi^{i}(u),\ \ \ \ y^{i}=\phi^{i}_{a}v^{a},\ \ \ \ \phi^{i}_{a}=\frac{\partial\phi^{i}}{\partial u^{a}}.$ (2.8)
For simplicity, we will denote $\phi(x)$ by $x$ and $d\phi X$ by $X$. Define
$\mathcal{V}(M)=\{(x,\xi)~|~x\in M,\xi\in T_{x}^{*}N,\xi(X)=0,\forall X\in T_{x}M\},$
which is called the normal bundle of $M$ [18]. Set $\tau\in\mathcal{N}M={\mathcal{L}}^{-1}(\mathcal{V}(M))\subset TN$, which satisfies $g_{\tau}(\tau,X)=0$ for any $X\in TM$. $\tau$ is called a normal vector field of $M$, and the locally Riemannian manifold $(M,\hat{g}=\phi^{*}g_{\tau})$ is called an anisotropic submanifold of $(N,F)$. Let $\textbf{n}\in\mathcal{N}^{0}M={\mathcal{L}}^{-1}(\mathcal{V}^{0}(M))$ be the unit normal vector field, where $\mathcal{V}^{0}(M)=\{\nu\in\mathcal{V}(M)|F^{*}(\nu)=1\}$ is the unit normal bundle. In general, $\mathcal{L}^{-1}(-\nu)\neq-\mathcal{L}^{-1}(\nu)$. The shape operator  ${A}_{\textbf{n}}:T_{x}M\rightarrow T_{x}M$ is defined by
${A}_{\textbf{n}}X=-(\nabla^{\textbf{n}}_{X}\textbf{n})^{\top}.$ (2.9)
We call the eigenvalues of ${A}_{\textbf{n}}$, $k_{1},k_{2},\cdots,k_{m}$, the principal curvatures of $M$ with respect to n and $\hat{H}_{\textbf{n}}=\sum\limits_{a=1}^{m}k_{a}$ the anisotropic mean curvature of $M$.
[…]
${A}_{\textbf{n}}X=-\nabla^{\textbf{n}}_{X}\textbf{n}.$ (2.10)
Let $\rho$ be a $C^{\infty}$ distance function on an open subset $U\subset(N,F)$ and $M_{s}:=\rho^{-1}(s)\subset U$ such that $M_{0}=M\cap U\neq\varnothing$. Let $\gamma(s)$ be a unit speed normal geodesic of the regular level hypersurface $M_{s}$. Denote $T(s)=\nabla\rho\in\mathcal{N}^{0}M_{s}$ and $A_{T(s)}$ as the shape operator of $M_{s}$. For $X\in TM_{s}$, define
$\dot{A}_{T(s)}X=\nabla^{T(s)}_{T(s)}(A_{T(s)}X)-A_{T(s)}(\nabla^{T(s)}_{T(s)}X).$
From [17], we have
$\dot{A}_{T(s)}=A_{T(s)}^{2}+\textbf{R}_{T(s)}.$ (2.11)
(2.11) is called the Riccati equation.
2.3 Isoparametric functions and isoparametric hypersurfaces
More precisely, the function $f$ is called isoparametric (resp. $d\mu$-isoparametric) on $(N,F,d\mu)$, if (1.2) holds for $\Delta f=\hat{\Delta}f$ (resp. $\Delta f=\Delta_{\sigma}f$), which is defined by (2.6) (resp. (2.7)). If $(N,F,d\mu)$ has constant $\mathbf{S}$-curvature, then $f$ is isoparametric if and only if it is $d\mu$-isoparametric. For simplicity, we only consider $\Delta f$ as $\hat{\Delta}f$ in this paper. Obviously, all conclusions in this paper still hold for $d\mu$-isoparametric hypersurfaces if $\mathbf{S}$-curvature is a constant. Similar to Riemannian isoparametric hypersurfaces, we can give a geometric definition of isoparametric functions by anisotropic mean curvatures and principal curvatures.
Definition 2.1.
[9] A transnormal function $f$ is isoparametric on $(N,F)$ if and only if each regular level hypersurface of $f$ has constant anisotropic mean curvatures.
Definition 2.2.
[9] Let $(N,F)$ be a Finsler manifold with constant flag curvature. Then a transnormal function $f$ is isoparametric if and only if each regular level hypersurface of $f$ has constant principal curvatures.
Unlike Riemannian case, there exists some locally isoparametric hypersurfaces that cannot be extended to global isoparametric hypersurfaces (See Remark 4.3). Indeed, if some conditions are added, such extension can be achieved (See Remark 5.1). Hence, we need to give the definition of locally isoparametric functions on $(N,F)$. If there exist a neighborhood $V$ of $M$ in $(N,F)$ and an isoparametric function $f$ defined on $V$ such that $M$ is a regular level hypersurface of $f$, then $f$ is called a locally isoparametric function and $M$ is called a locally isoparametric hypersurface[13]. We can also give the geometric definition of locally isoparametric hypersurfaces in $(N,F)$.
Definition 2.3.
[7] A hypersurface $M$ in a Finsler manifold is a locally isoparametric hypersurface if and only if it and its sufficiently close parallel hypersurfaces have constant anisotropic mean curvatures.
Definition 2.4.
[13] $M$ is a locally isoparametric hypersurface in a Finsler manifold with constant flag curvature if and only if its principal curvatures are all constants.
3 Transnormal functions on Finsler space forms
3.1 Transnormal functions and focal varieties
For $\tau\in\mathcal{N}_{x}M$, define the normal exponential map $E:\mathcal{N}M\rightarrow N$ satisfying $E(x,\tau)=\exp_{x}\tau$, which is $C^{\infty}$ on $TN\setminus\{0\}$. The focal points of $M$ are the critical values of the normal exponential map $E$. The focal point $p\in N$ has multiplicity $m$ means that $E_{*}$ at $(x,\tau)$ has nullity $m>0$. The set of focal points is called a focal set, which is denoted by $\textmd{Foc}(f)$. The focal varieties of $f$ is denoted by $M_{\pm}$, which is the set where $f$ attains its global maximum value or global minimum value (if either of them exists). Denote $V=\{p\in N\big{|}\nabla f(p)=0\}$, which is the set of critical point of $f$. Indeed, $\textmd{Foc}(f)\subset M_{\pm}\subset V$.
Lemma 3.1.
[5] Let $f$ be a transnormal function on $(N,F)$. Set $\alpha,\beta]\subset\overline{J}$. If $x\notin f(V)$, where $x\in[\alpha,\beta)$, and $\beta\in f(V)$, then
$d(M_{\alpha},M_{\beta})=\int^{\beta}_{\alpha}\frac{df}{\sqrt{a(f)}}$ (3.1)
and the improper integral in (3.1) converges.
Remark 3.2.
In fact, from the proof of Lemma 3.1 in [5], we know that if $x\notin f(V)$, where $x\in(\alpha,\beta)$, and $\alpha,~\beta\in f(V)$, then (3.1) still holds.
Define the reverse metric of $F$ by $\overleftarrow{F}(x,y)=F(x,-y)$, where $y\in T_{x}N$, $x\in N$. Then we have
Lemma 3.3.
[5] Let $(N,F)$ be a complete, connected Finsler manifold and $f$ be an appropriate transnormal function on $N$, then we have
(1)each focal variety of $f$ is a smooth submanifold in $N$.
(2)each regular level set of $f$ is a tube or a parallel hypersurface over either of the focal varieties with respect to $\overleftarrow{F}$ or $F$.
[…]
Lemma 3.4.
Let $f$ be an appropriate transnormal function on $(N,F)$. If $\gamma(s)$ is a maximal $f$-segment satisfying $\gamma(0)\in M_{+}$ (resp. $M_{-}$), then $\gamma^{\prime}(0)\in\mathcal{N}^{0}_{\gamma(0)}M_{+}$ (resp. $\mathcal{N}^{0}_{\gamma(0)}M_{-}$). Conversely, for any $\eta\in\mathcal{N}^{0}_{p}M_{+}$ (resp. $\mathcal{N}^{0}_{q}M_{-}$), there exists a maximal $f$-segment $\gamma(s)$ satisfying $\gamma(0)=p\in M_{+}$ (resp. $\gamma(0)=q\in M_{-}$) and $\gamma^{\prime}(0)=\eta$.
Remark 3.5.
If $f$ is only a transnormal function and $M_{\pm}$ are submanifolds, Lemma 3.4 holds.
For an appropriate transnormal function $f$ satisfying $F^{2}(\nabla f)=a(f)$, from Lemma 3.1, we define $d_{+}=d(M_{-},M_{+})=\int^{\beta}_{\alpha}\frac{dt}{\sqrt{a(t)}}$. Let f-f be a transnormal function satisfying $F^{2}(\nabla(-f))=\tilde{a}(-f)$. Due to $\nabla f=0$ if and only if $\nabla(-f)=0$, we know that f-f has no critical values in $f(N_{f})$. Since $M_{\pm}^{f}=M_{\mp}^{-f}$, we define $d_{-}=d(M_{+},M_{-})=\int^{\beta}_{\alpha}\frac{dt}{\sqrt{\tilde{a}(t)}}$. From Lemma 3.4, we immediately have
Lemma 3.6.
Let $f$ be an appropriate transnormal function and f-f be a transnormal function on $(N,F)$. If $d_{+}<+\infty$ and $\gamma(s)$, $s\in[0,+\infty)$, is a unit speed geodesic such that $\gamma(s)\big{|}_{[0,d_{+}]}$ is a maximal $f$-segment, then $\gamma(s)\big{|}_{[d_{+},d_{+}+d_{-}]}$ is a maximal (f)(-f)-segment with respect to $F$.
Remark 3.7.
If $f$ is a transnormal function with no critical values in $J$ and $M_{\pm}$ are submanifolds, then Lemma 3.6 holds.
3.2 Proof of Theorem 1.1
Lemma 3.8.
[12] Let $\phi:M\rightarrow\overrightarrow{N}(c)$ be an immersion submanifold and $\mathbf{n}$ be a unit normal vector to $\phi(M)$ at $x$. Then $p=\phi_{s}(x,\mathbf{n})$ is a focal point of $(M,x)$ of multiplicity $m_{0}>0$ if and only if there is an eigenvalue $\kappa$ of the shape operator $A_{\mathbf{n}}$ of multiplicity $m_{0}$ such that
$\kappa(x)=\left\{\begin{array}[]{lcl}\frac{1}{s},&&{c=0,}\\ \sqrt{c}\cot\sqrt{c}s,&&{c>0,}\\ \sqrt{-c}\coth\sqrt{-c}s,&&{c<0.}\end{array}\right$. (3.5)
Let $M=f^{-1}(t_{0})$ be an oriented regular hypersurface by $\nabla f$ and $\kappa_{1}$, $\kappa_{2}$, $\ldots$, $\kappa_{g}$ be its distinct principal curvatures at $x$. Let $\gamma(s):(-\infty,\infty)\rightarrow N$, be a geodesic satisfying $\gamma(0)=x$ and $\gamma(r)=p\in M_{+}$, where $r=d(M,M_{+})=\int^{\beta}_{t_{0}}\frac{df}{\sqrt{a(f)}}$ by Lemma 3.1. Then
Case 1.
$c>0$.
In this case, $d_{+},d_{-}<+\infty$. From Lemma 3.8, for any $\kappa_{i}$, there exists some $s_{i}$ such that $\gamma(s_{i})\in Foc(f)$. Then From Lemma 3.6 and (3.5), there must exist $n_{i},m_{i}\in Z$ such that the principal curvatures can be expressed by
$\kappa_{i}=\sqrt{c}\cot\sqrt{c}(r+n_{i}d_{-}+m_{i}d_{+}).$
Case 2.
$c=0$.
From Lemma 3.8, for any $\kappa_{i}\neq 0$, there exists some $s_{i}$ such that $\gamma(s_{i})\in Foc(f)$.
(1) $Foc(f)=\varnothing$. Obviously, $g=1$ and $\kappa_{1}=0$. Then we must have $\kappa_{1}\equiv 0$ since otherwise $Foc(f)\neq\varnothing$.
(2) $Foc(f)\neq\varnothing$. If $Foc(f)=M_{+}$, then $\kappa_{1}=\frac{1}{r}$. If $Foc(f)=M_{-}$ and $Foc(f)\neq M_{+}$, then $\kappa_{1}=\frac{1}{r+d_{-}}$. From [5], $p=\gamma(r_{0})$ is the focal point defined by the normal exponential map $\Phi(s,t):=E(\sigma(t),s\textbf{n}(\sigma(t)))$, where $\sigma(t)$ is a curve on $M$ and $\textbf{n}(\sigma(t))\in\mathcal{N}^{0}_{\sigma(t)}M$, if and only if $J(r_{0})=\frac{\partial\Phi}{\partial t}\big{|}_{t=0,s=r_{0}}=0$. From Cartan-Hadamard Theorem, $\gamma(s)$ passes through $Foc(f)$ once. Hence, from Lemma 3.6 and (3.5), we have $g=2$ and $\kappa_{2}=0$.
Case 3.
$c<0$.
From Lemma 3.8, for any $|\kappa_{i}|>\sqrt{-c}$, there exists $s_{i}$ such that $\gamma(s_{i})\in Foc(f)$. For all $|\kappa_{i}|\leq\sqrt{-c}$, $Foc(f)=\varnothing$.
From Cartan-Hadamard Theorem, there exists one $s_{i}$ at most such that $|\kappa_{i}|>\sqrt{-c}$. Similar to the analysis in Case 2, from Lemma 3.6 and (3.5), if $|\kappa_{i}|\geq\sqrt{-c}$, then the principal curvatures may be $\sqrt{-c}\coth\sqrt{-c}r$, $\sqrt{-c}\coth\sqrt{-c}(r+d_{-})$, $\sqrt{-c}$ or $\sqrt{-c}$, and $g\leq 2$.
Whether it is case 1, case 2 and case 3, the principal curvatures $\kappa_{i}$ of $M$ are all constants. From Definition 2.4, $M=f^{-1}(t_{0})$ is a locally isoparametric hypersurface and $f$ is a locally isoparametric function. Further more, from (2.11),
$-\kappa^{\prime}_{i}(s)=\kappa_{i}^{2}(s)+c,$ (3.6)
which is an ODE with respect to an arc-length parameter $s$ and its initial values are determined by the constant principal curvatures of $M$. Combine Lemma 3.6 and (3.6), we know each level hypersurface of $f$ has constant principal curvatures. From Definition 2.2, $M=f^{-1}(t_{0})$ is a global isoparametric hypersurface and $f$ is a global isoparametric function. This completes the proof of Theorem 1.1. From the proof of Theorem 1.1, we immediately have
Theorem 3.9.
Let $(N(c),F)$ be a Finsler space form. Suppose that $f$ is an appropriate transnormal function and f-f is a transnormal function on $N(c)$. If $c=0$, or $c<0$ and all principal curvatures $\kappa_{i}$ of a regular level hypersurface satisfying $|\kappa_{i}|\geq\sqrt{-c}$, $1\leq i\leq g$, then each regular level hypersurface has two distinct principal curvatures at most.
Remark 3.10.
In [12], we obtain the same result in different conditions. In fact, we need to rectify an omission of Theorems in [12]. That is, for $c<0$ and $c=0$, it also need to satisfy $|\kappa_{i}|>\sqrt{-c}$ and $\kappa_{i}\neq 0$, respectively.
Remark 3.11.
In local case, Theorem 3.9 doesn’t hold. That is, there exist locally isoparametric hypersurfaces in some Minkowski spaces with more than two principal curvatures (See Theorem 4.2 for detail).
If $F$ is reversible, $\tilde{a}(-f)=F^{*}(-df)=F^{*}(df)=F(\nabla f)=a(f)$. That is, $f$ is an appropriate transnormal function if and only if f-f is an appropriate transnormal function. Hence, we have
Corollary 3.12.
Let $(N(c),F)$ be a Finsler space form with a reversible metric $F$ and $f$ be an appropriate transnormal function on $N(c)$. Then $f$ is isoparametric when $c\geq 0$, or $c<0$ and all principal curvatures $\kappa_{i}$ of a regular level hypersurface satisfying $|\kappa_{i}|\geq\sqrt{-c}$, $1\leq i\leq g$.
4 Locally isoparametric hypersurfaces
4.1 The number of distinct principal curvatures
Theorem 1.1 and Theorem 3.9 are not necessarily true for locally transnormal functions. For example, the distance function to a hypersurface is a locally appropriate transnormal function for a reversible Finsler metric. But it is not necessarily isoparametric. Even for locally isoparametric functions, The conclusion of Theorem 3.9 is not necessarily true.
In [13], the conic Finsler metric was defined and a local minimal isoparametric hypersurface was finded, which are not conic Minkowski hyperplanes, conic Minkowski hyperspheres or conic Minkowski cylinders.
Lemma 4.1.
[13] Let $(\mathbf{R}^{3},F)$ be a 3-dimensional conic Minkowski-$(\alpha,\beta)$ space and the dual metric of $F$ be $F^{*}=\alpha^{*}\varphi(\frac{\beta^{*}}{\alpha^{*}})$, where
$\varphi=\frac{\sqrt{b^{2}-(1+a^{2})s^{2}}}{b}-\frac{as}{b}\arctan\frac{\sqrt{b^{2}-(1+a^{2})s^{2}}}{as},0<|s|<\frac{b}{\sqrt{1+a^{2}}},$
where $\beta^{*}=(0,0,b)$ and $a,b$ are two positive constants. Then in $(\mathbf{R}^{3},F)$, helicoid $\mathbf{r}=(u\cos v,u\sin v,av)(0<u<1)$ is a local minimal isoparametric hypersurface with constant principal curvatures $\pm 1$.
In fact, we can construct a regular Finsler metric $\tilde{F}$ on $\mathbf{R}^{n}$ by the conic Finsler metric constructed in Lemma 4.1. Set $x_{0}=\frac{b}{(\phi-s\phi^{\prime})\sqrt{b^{2}-s^{2}}}$ and $x_{1}=x_{0}-\frac{1}{2a(\phi-s\phi^{\prime})\sqrt{b^{2}-s^{2}}}(s(\phi-s\phi^{\prime})+b^{2}\phi^{\prime})>x_{0}$. Denote
$L:\{\xi=(\xi_{1},0,\xi_{3})|F^{*}(\xi)=1,x_{0}<\xi_{1}<x_{1}\}.$
Then we obtain a strictly convex surface $\Sigma$ by rotating $L$, which is in a conic domin
$U=\{\xi\in\mathbf{R}^{3}~\big{|}~\xi_{1}^{2}+\xi_{2}^{2}>4a^{2}\xi_{3}^{2},\xi_{3}^{2}>0\}.$
Since $\frac{d\xi_{3}}{d\xi_{1}}\big{|}_{x_{0}}>0$, we can obtain a strictly convex smooth and closed surface $\tilde{\Sigma}$ including the origin in $\mathbf{R}^{3}$ by extending $\Sigma$.
We can define a regular Minkowski metric $\overline{F}^{*}$ satisfying $\overline{F}^{*}\mid_{\Sigma}=F^{*}$ on $\mathbf{R}^{3}$. Let $\overline{F}$ be the dual metric of $\overline{F}^{*}$. Obviously, $\overline{F}$ is still a regular Minkowski metric. In Lemma 4.1, the normal vector of the helicoid r is in $U$ if and only if $0<u<\frac{1}{2}$. Hence, $\textbf{r}=(u\cos v,u\sin v,av)$ ($0<u<\frac{1}{2}$) is a locally isoparametric hypersurface in $(\mathbf{R}^{3},\overline{F})$ with constant principal curvature $\pm 1$.
[…]
Theorem 4.2.
In some Minkowski spaces, there exist locally isoparametric hypersurfaces with three distinct principal curvatures at least.
From the analysis process of Theorem 4.2, we also know that
Remark 4.3.
There exist locally minimal isoparametric hypersurfaces in some Minkowski spaces, which are not Minkowski hyperplanes, Minkowski hyperspheres or Minkowski cylinders. Hence, locally isoparametric hypersurfaces may not be extended to global isoparametric hypersurfaces in Finsler manifolds, which is different from Riemannian case.
4.2 Umbilic hypersurfaces
In [9], we know that any totally umbilic hypersurface in a Minkowski space must be isoparametric. In this section, we generalize to the following,
Theorem 4.4.
Let $(N,F)$ be a Finsler space with constant flag curvature. Then a totally umbilic hypersurface is locally isoparametric.
Proof.
Suppose that $M$ is a totally umbilic hypersurface of $N$ with principal curvature $\lambda$, from (2.8) and (2.10),
$D^{\textbf{n}}_{\partial u_{a}}\textbf{n}=-\lambda d\phi(\partial u_{a}).$ (4.1)
Set $\textbf{n}=n^{i}\frac{\partial}{\partial x^{i}}$, $n^{i}_{a}=\frac{\partial n^{i}}{\partial u^{a}}$ and $n^{i}_{ab}=\frac{\partial n^{i}}{\partial u^{a}\partial u^{b}}$. From (2.3),
$D^{\textbf{n}}_{\partial u_{b}}D^{\textbf{n}}_{\partial u_{a}}\textbf{n}$ $=D^{\textbf{n}}_{\partial u_{b}}(n^{j}_{a}+n^{i}\phi^{l}_{a}\Gamma^{j}_{li}(\textbf{n}))\partial x_{j}=\Big{(}n^{j}_{ab}+n^{i}\phi^{l}_{ab}\Gamma^{j}_{li}(\textbf{n})+(n^{i}_{b}\phi^{l}_{a}+n^{l}_{a}\phi^{i}_{b})\Gamma^{j}_{li}(\textbf{n})$
$+n^{i}\phi^{l}_{a}\phi^{p}_{b}\big{(}\frac{\partial\Gamma^{j}_{li}}{\partial x^{p}}(\textbf{n})+\frac{\partial\Gamma^{j}_{il}}{\partial y^{q}}(\textbf{n})\frac{\partial n^{q}}{\partial x^{p}}+\Gamma^{k}_{li}(\textbf{n})\Gamma^{j}_{pk}(\textbf{n})\big{)}\Big{)}\partial x_{j}.$ (4.2)
From (2.2),
$n^{i}\phi^{l}_{a}\phi^{p}_{b}(\frac{\partial\Gamma^{j}_{li}}{\partial x^{p}}(\textbf{n})+\frac{\partial\Gamma^{j}_{il}}{\partial y^{q}}(\textbf{n})\frac{\partial n^{q}}{\partial x^{p}})=n^{i}\phi^{l}_{a}\phi^{p}_{b}\frac{\delta\Gamma^{j}_{li}}{\delta x^{p}}(\textbf{n}).$
Let $c$ be the constant flag curvature of $N$, then
$R^{j}_{ipl}=c(g_{il}\delta^{j}_{p}-g_{ip}\delta^{j}_{l}).$
Hence, from (4.2),
$D^{\textbf{n}}_{\partial u_{b}}D^{\textbf{n}}_{\partial u_{a}}\textbf{n}-D^{\textbf{n}}_{\partial u_{a}}D^{\textbf{n}}_{\partial u_{b}}\textbf{n}$ $=n^{i}\phi^{l}_{a}\phi^{p}_{b}R^{j}_{ipl}(\textbf{n})=cn^{i}(\phi^{j}_{b}\phi^{l}_{a}g_{il}(\textbf{n})-\phi^{p}_{b}\phi^{l}_{a}g_{ip}(\textbf{n}))$
$=c\phi^{j}_{b}g_{\textbf{n}}(\textbf{n},d\phi(\partial u_{a}))-c\phi^{j}_{a}g_{\textbf{n}}(\textbf{n},d\phi(\partial u_{b}))=0.$ (4.3)
From (4.1) and (4.2), by a direct calculation,
$0=D^{\textbf{n}}_{\partial u_{b}}(-\lambda d\phi(\partial u_{a}))-D^{\textbf{n}}_{\partial u_{a}}(-\lambda d\phi(\partial u_{b}))=\frac{\partial\lambda}{\partial u^{a}}\phi^{j}_{b}-\frac{\partial\lambda}{\partial u^{b}}\phi^{j}_{a}.$
Hence, for any $a\neq b$, $\lambda$ is a constant. From Definition 2.4, $M$ is a locally isoparametric hypersurface. This completes the proof. ∎
5 Proof of Theorem 1.3
Let $f$ be a transnormal function on a connected forward Finsler space form $(N(c),F)$ and $M$ be a totally umbilic level hypersurface of $f$. Since $M$ is totally umbilic, from Theorem 4.4, the principal curvatures of $M$ are constants. The solution of (3.6) is determined by the principal curvatures of $M$ and the distance (3.1) between hypersurfaces. Hence, for any $t\in J$, $f^{-1}(t)$ has constant principal curvatures. Combine Definition 2.2, $f$ is a global isoparametric function on $(N(c),F)$.
Remark 5.1.
From the proof of Theorem 1.3, we know that if $f$ is a transnormal function on a forward Finsler space form, then the locally isoparametric hypersurfaces can be extended to the global isoparametric hypersurfaces.
6 Proof of Theorem 1.4
From [9], it is easy to get that
Lemma 6.1.
Let $f$ be an isoparametric function on a Finsler manifold $(N,F)$ with constant flag curvature $c$. Then the principal curvature $\kappa_{i}$ of an isoparametric hypersurface $f^{-1}(t)$ satisfying
$\sum_{i=1}^{n-1}\kappa_{i}$ $=\frac{a^{\prime}(t)}{2\sqrt{a(t)}}-\frac{b(t)}{\sqrt{a(t)}},$ (6.1)
$\sqrt{a(t)}\frac{\partial\kappa_{i}}{\partial t}$ $=c+\kappa_{i}^{2},\quad i=1,\cdots,n-1.$ (6.2)
Let $(N(c),F)$ be a forward Finsler space form and $r(x)=d_{F}(p,x)$ be the distance function from $p$ to $x$. Set $f=\frac{1}{2}r^{2}$. Obviously, $f$ is a non-constant $C^{1}$ function on $\mathcal{D}_{p}$, which is smooth on $\mathcal{D}_{p}\setminus\{p\}$.
By a direct calculation, $F(\nabla f)=\sqrt{2f}$. From the proof of Theorem 4.1 in [16], we have $\hat{\Delta}r=(n-1)\frac{\mathfrak{s}^{\prime}_{c}(r)}{\mathfrak{s}_{c}(r)}$, where
$\mathfrak{s}_{c}(r)=\left\{\begin{array}[]{lcl}r,&&{c=0,}\\ \frac{\sin\sqrt{c}r}{\sqrt{c}},&&{c>0,}\\ \frac{\sinh\sqrt{-c}r}{\sqrt{-c}},&&{c<0.}\end{array}\right$. (6.6)
Then
$\hat{\Delta}f=r\hat{\Delta}r+1=r(n-1)\frac{\mathfrak{s}^{\prime}_{c}(r)}{\mathfrak{s}_{c}(r)}+1=\sqrt{2f}(n-1)\frac{\mathfrak{s}^{\prime}_{c}(\sqrt{2f})}{\mathfrak{s}_{c}(\sqrt{2f})}+1.$
Hence, $f$ satisfies (1.2) with $a(t)=2t,b(t)=\sqrt{2t}(n-1)\frac{\mathfrak{s}^{\prime}_{c}(\sqrt{2t})}{\mathfrak{s}_{c}(\sqrt{2t})}+1$. Therefore, $f$ is a locally isoparametric function on $\mathcal{D}_{p}$.
By (6.1) and (6.2), a straightforward calculation yields that
$\sum\limits_{a}\kappa_{a}=\left\{\begin{array}[]{lcl}\frac{1}{\sqrt{2t}}(1-n),&&{c=0,}\\ -(n-1)\sqrt{c}\cot(\sqrt{2ct}),&&{c>0,}\\ -(n-1)\sqrt{-c}\coth(\sqrt{-2ct}),&&{c<0,}\end{array}\right$. (6.10)
and
$\sum\limits_{a}\kappa_{a}^{2}=\left\{\begin{array}[]{lcl}\frac{1}{2t}(n-1),&&{c=0,}\\ (n-1)c(\csc^{2}(\sqrt{2ct})-1),&&{c>0,}\\ -c(n-1)(\textmd{csch}^{2}(\sqrt{-2ct})+1),&&{c<0.}\end{array}\right$. (6.14)
(6.10) and (6.14) imply that
$(\sum_{a}\kappa_{a})^{2}$ $=(n-1)\sum_{a}\kappa^{2}_{a}.$ (6.15)
From (6.10) and (6.15), we have
$\kappa_{1}=\kappa_{2}=\ldots=\kappa_{n-1}=\left\{\begin{array}[]{lcl}-\frac{1}{\sqrt{2t}},&&{c=0,}\\ -\sqrt{c}\cot(\sqrt{2ct}),&&{c>0,}\\ -\sqrt{-c}\coth(\sqrt{-2ct}),&&{c<0.}\end{array}\right$.
That is, for a fixed level hypersurface $M_{t}$, $\kappa_{a}(t)$ is a constant.
Let $\mathbb{S}^{n}$ be an $n$-sphere equipped with a Finsler metric $F$ and the Busemann-Hausdorff volume form $d\mu$. If it has constant flag curvature $c$, vanishing S-curvature and $Diam=\frac{\pi}{\sqrt{c}}$, we call it a standard Finsler sphere and denote it by $(\mathbb{S}^{n}(c),F,d\mu)$[19]. We can obtain a global isoparametric function on $(\mathbb{S}^{n}(c),F,d\mu)$ with $g=1$. Set $f(x)=-\cos(\sqrt{c}r(x))$. From [19], we know that $f(x)$ is the first eigenfunction of $(\mathbb{S}^{n}(c),F,d\mu)$. Hence, from [8], $f\in C^{\infty}(\mathbb{S}^{n}_{f})$ and $f\in C^{1}(\mathbb{S}^{n})$. In addition, since $F(\nabla f)=\sqrt{c(1-f^{2})}$ and $\Delta f=-ncf$, then $f$ is isoparametric on $(\mathbb{S}^{n}(c),F,d\mu)$ with $a(t)=c(1-t^{2})$ and b(t)=nctb(t)=-nct. From Lemma 6.1, we also have $\kappa_{1}=\kappa_{2}=\cdots=\kappa_{n-1}=\frac{\sqrt{c}t}{\sqrt{1-t^{2}}}$. That completes the proof of Theorem 1.4.